OpenStudy (anonymous):

Simplify 66sqrt 77 divided by 11 sqrt 7

OpenStudy (akshay_budhkar):

\[{66\sqrt{77} \over { 11 \sqrt{7} } }=6\sqrt{11}\]

OpenStudy (anonymous):